OpenStudy (anonymous):

Find all the possible rational roots of: 1. f(x)=2x^4-5x^3+8x^2+4x+7 2. f(x)=x^3+4x^2-3x-5 3. f(x)=x^3+25

OpenStudy (anonymous):

1. Step 1 : Simplify 2x4-5x3+8x2+4x + 7 Polynomial Roots Calculator : 1.1 Find roots (zeroes) of F(x) = 2x4-5x3+8x2+4x+7 Polynomial Roots Calculator is a set of methods aimed at finding values of x for which F(x)=0 Rational Roots Test is one of the above mentioned tools. It would only find Rational Roots that is numbers x which can be expressed as the quotient of two integers The Rational Root Theorem states that if a polynomial zeroes for a rational number P/Q then P is a factor of the Trailing Constant and Q is a factor of the Leading Coefficient In this case, the Leading Coefficient is 2 and the Trailing Constant is 7. The factor(s) are: of the Leading Coefficient : 1,2 of the Trailing Constant : 1 ,7 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 18.00 -1 2 -0.50 7.75 -7 1 -7.00 6888.00 -7 2 -3.50 605.50 1 1 1.00 16.00 1 2 0.50 10.50 7 1 7.00 3514.00 7 2 3.50 204.75 Polynomial Roots Calculator found no rational roots Equation at the end of step 1 : 2x4 - 5x3 + 8x2 + 4x + 7 = 0 Step 2 : Solve 2x4-5x3+8x2+4x+7 = 0 Quartic Equations : 2.1 Solve 2x4-5x3+8x2+4x+7 = 0 In search of an interavl at which the above polynomial changes sign, from negative to positive or the other wayaround. Method of search: Calculate polynomial values for all integer points between x=-20 and x=+20 No interval at which a change of sign occures has been found. Consequently, Bisection Approximation can not be used. As this is a polynomial of an even degree it may not even have any real (as opposed to imaginary) roots No Solutions found